OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Enter your answer as an ordered pair. 3x+y=10 y=x-2

OpenStudy (anonymous):

My favourite. K so what you do is very simple. you have two equation, label them as (1) and (2). 3x+y=10 (1) y = x+ 2 (2) Substitute (2) into (1) Try it.

OpenStudy (anonymous):

Do you know how to substitute?

OpenStudy (anonymous):

no i suck at algerbra

OpenStudy (anonymous):

substitute 2 into 1; 3x+x-2=10 4x=12 x=3

OpenStudy (anonymous):

alright so this is what you do. 3x+y = 10 (1) y = x+2 (2) Substitute (2) into (1) 3x + (x+2) = 10 4x + 2 = 10 4x = 10 -2 4x = 8 *Divide both sides by 4* x = 2

OpenStudy (anonymous):

that is your x value. Now for your y value you do the same thing. Substitute x = 2 into (1) 3x + y = 10 3(2) + y = 10 6 + y = 10 *Subtract 6 from both sides* y = 4

OpenStudy (anonymous):

Therefore your ordered pair is (2,4). This is the point of intersection :)

OpenStudy (anonymous):

@OrthodoxMan y=x-2 :)

OpenStudy (anonymous):

@josiahh thanks man i cant believe i couldnt read that

OpenStudy (anonymous):

well then Josiah is right about the x value. Substitute x = 3 into (1) and you get 3x + y = 10 3(3) = y = 10 y = 10 - 9 y = 1 Therefore the ordered pair is (3,1)